# Identifying Materials-Level Sources of Performance Variation in Superconducting Transmon Qubits

### v1

@@ II   Results
In the ensuing sections, we focus solely on the variety of structural, chemical, and physical differences from qubit-to-qubit within a single chip using these techniques. We first focus on the few parameters that appeared to correlate strongly with qubit-to-qubit variations in the average quality factor and include electromagnetic (EM) simulations as needed to probe the origin of these trends. The quality factor was selected as the primary performance metric of interest to account for the spread in frequency from qubit-to-qubit on a single chip. We then discuss potential proxy techniques for predicting qubit performance, as well as observed variations from device to device that do not exhibit a trend with performance.
II.1Relationship between Trench Depth and Performance
One of the major observations from this study was the trend between variation in qubit performance and variations in trench depth as demonstrated in Figure 2. In agreement with previous reports by Gambetta et al. and Murray, we observe that there appears to be a relationship between increasing trench depth, which represents the depth of recesses etched into the underlying dielectric substrate immediately adjacent to superconducting metal edges, and increased qubit quality factor within qubits on a single chip [24, 18, 25]. An example of the relationship between trench depth and the quality factor of the qubit is provided in Figure S4 for superconducting qubits prepared with a niobium base metal capped with tantalum metal [1] on a silicon substrate.
@@
Based on electromagnetic (EM) simulations of the qubit geometry employed in this study, we observe that the energy participation ratio (EPR) of the lossy surface dielectric decreases logarithmically with increases in trench depth and saturates at trench depth values of the order of 70 nm, as seen in Figure 2b. We find that this reduction in the surface EPR primarily arises from a suppression in the field at the sidewalls (side EPR) of the superconducting metallization layer with increasing trench depth as opposed to suppression in the field from the top surface (top EPR). Qualitatively, these simulations agree with the experimental results regardless of the superconducting materials and substrates used, where, similarly, small variations in the trench depth trend with significant variations in the qubit quality factor for qubits with trench depth values less than 20 nm. With increasing trench depth, variations in trench depth appear to induce smaller variations in the qubit quality factor, and in the case where trench depths exceed 70 nm, trench depth appears to be uncorrelated with the qubit quality factor. However, these variations in trench depth do not entirely explain the qubit-to-qubit variations in performance observed in this study. Based on the simulated data, the shaded regions in 2a represent the level of qubit-to-qubit variation for a single chip expected in the case where trench depth represents the sole source of variation. This assumes a 10% error in the loss predicted by the simulations. Given that the performance of several qubits lies outside the shaded regions, it appears that other sources of variation contribute to the variations in performance observed. In the next sections, we will discuss two additional sources that correlate with performance. These sources are most apparent for qubits with substrate trench depths exceeding 70 nm (Chips 4 and 6) and qubits with a top surface consisting of niobium oxide (Chip 2).
Figure 4: EM simulations of the qubit sidewall. The sidewall angle ($\alpha$), oxide thickness (dh), radius of curvature of the top edge (R1), radius of curvature of the bottom edge (R2), distance of undercut ($x$), and undercut angle ($\beta$) are defined in (a). Surface EPR values are provided as a function of (b) $\alpha$ for dh = 5 nm and (c) dh for $\alpha$ = 0 degrees. Top represents the EPR associated with the top surface, side represents the EPR associated with the sidwall surface, and sum represents the total surface EPR, i.e. the sum of these two quantities.
II.2Sidewall Geometry
@@
As seen in Figure 4, we find that increasing $\alpha$ leads to an increase in the surface EPR of the sidewall. The simulations indicate that devices with $\alpha$ of approximately 30 degrees (as seen in Figure 3) exhibit surface EPR values that are 15-20% larger than a comparable device with perpendicular sidewalls. On the other hand, we find that variations in the radius of curvature of the top edge have virtually no impact on the surface EPR due to competing effects between fields in the top surface and sidewall. Similarly, virtually no impact on surface EPR is observed when the radius of curvature of the base of the film varies, which also supports our experimental findings (Figure S6).
II.3Niobium oxide
@@
In particular, the data set associated with Chip 2 in Figure 2 represents a set of 4 qubits where the base metal consisted of bare niobium. As seen in Figure 5, although there is no correlation between trench depth and variations in qubit quality factor between devices, increases in niobium oxide thickness, specifically at the top surface, trend with decreases in quality factor. This agrees with the simulations in Figure 4. The bars in the scatter plot represent the range in thicknesses observed for each device and indicate that the surface oxide thickness of this polycrystalline film of niobium varies from 5 nm to 7 nm across a single 7.5 mm by 7.5 mm chip. The shaded regions in 5e represent the level of qubit-to-qubit variation expected in the case where the thickness of the niobium oxide represents the sole source of variation based on the simulated data.
@@
II.5Inconclusive Impact on Performance
@@
II.5Macroscopic Debris and Lithographic Defects
@@
II.5Josephson Junction: Structure and Electrical Properties
@@
As illustrated in Figures S9 and S10, variations in grain structure and sidewall geometry of the Josephson Junction electrodes, as well as I-V response, are observed for the measured qubit devices. Concerning grain structure, we observe that the median grain size in the upper aluminum lead can vary from 30 nm to 50 nm, depending on the qubit. Similarly, in the lower aluminum lead, we observe that the median grain size varies from 20 nm to 35 nm. In both cases, there does not appear to be any relationship with the qubit quality factor. Similar variations to those observed in the sidewall angle of superconducting pads are observed for the aluminum leads, as captured in Figure S9. In this case, however, no clear trends between sidewall angle and qubit performance are observed. Finally, in terms of the I-V response, device-to-device variations are also observed in terms of normal state resistance as well as switching current (Figure S10). Once again, no trend is observed between these parameters and qubit performance, which thereby suggests that in the case of the high-coherence superconducting qubits investigated in this study, variations in the Josephson Junction morphology do not appreciably alter the overall performance.
II.5Surface Chemistry
@@
We employed ToF-SIMS to probe the surface chemistry of the superconducting pads. The resulting average impurity counts are provided in Figure S11 Across the three analyzed devices, there do not appear to be any obvious trends between the surface chemistry of the superconducting pads and the qubit performance. Potentially, there may be a trend between residual fluorine content and performance, however, increased statistics are needed to isolate the impact of this impurity on performance. Moreover, although hydrogen impurities are known to bind with niobium at cryogenic temperatures and form non-superconducting precipitates, as reported by Lee et al. and seen in Figure S12, no consistent trend is observed between hydrogen impurity content and performance across all qubits[38]. This is in agreement with recent work suggesting that hydrogen impurities in niobium may not be linked to TLS loss [39]. Additionally, we also probed the chemistry of the substrate surface adjacent to the measured qubits. In this case, the spot-to-spot variation in the substrate surface chemistry for a single qubit was found to be similar to the variation in substrate surface chemistry observed across devices. Thus, dedicated studies where the surface chemistry is locally altered are needed to understand the impact of impurities in superconducting pads and substrate regions on device performance.
II.5Sidewall Oxide Thickness
@@
Similar to variations in the niobium top surface oxide thickness observed in Figures 5, variations in the niobium sidewall oxide thickness are also observed and reported in these figures. Unlike our observations for the niobium surface oxide, there does not exist an obvious correlation between variations in niobium sidewall oxide thickness and variations in performance. This may be due to two factors. First, unlike the 1-2 nm variations observed in the niobium surface oxide thickness, the variations in the niobium sidewall oxide thickness are of a smaller magnitude and typically range from 0.5 nm -1 nm. Once again, an additional dedicated study focused on varying the thickness of the niobium oxide at the sidewall is needed to meaningfully extract the relative impact variations in this parameter imposed on performance.
@@ III   Discussion
Concerning sidewall geometries, perpendicular and reproducible sidewalls can be achieved by developing more anisotropic dry etching recipes for the metal. To this end, sidewall imaging with electron microscopy can be a valuable tool for recipe optimization, while parameters such as etching chemistry (F-based vs Cl-based), level of physical bombardment (Ar content), and power are varied. Finally, with respect to surface encapsulation [1], our results indicate that incomplete coverage, especially close to the sidewall, can have a noticeable detrimental effect on device performance. This was the case for niobium films capped with AuPd as opposed to niobium films capped with Ta.

### v2

@@ II   Results
In the ensuing sections, we focus primarily on the variety of structural, chemical, and physical differences from qubit-to-qubit within a single chip using these techniques as summarized in Figure 2. Upon discussing these observations, we then focus on the few parameters that appeared to correlate strongly with qubit-to-qubit variations in the average quality factor and include electromagnetic (EM) simulations as needed to probe the origin of these trends. The quality factor was selected as the primary performance metric of interest to account for the spread in frequency from qubit-to-qubit on a single chip. We finally discuss potential proxy techniques for predicting qubit performance, as well as observed variations from device to device that do not exhibit a trend with performance.
II.1Findings
As captured in Figure 2, the aforementioned characterization techniques reveal a bevy of structural, chemical, and physical differences from qubit-to-qubit within a single chip. OM and SEM analysis reveal the presence of macroscopic debris and lithographic defects in the vicinity of individual qubits as well as the neighboring readout resonators. S/TEM analysis reveals variations in the thickness and grain size of the Josephson junction leads. We do not observe meaningful differences in the chemistry or the thickness of the AlOx insulating barrier across qubits as analyzed with S/TEM-EELS. With respect to the superconducting Nb pads, this technique also reveals variations in the sidewall geometry, depth of the etched substrate trench, and oxide thicknesses of the top surface and the sidewall. We also do not observe meaningful differences in the chemistry of these oxides across a single chip.
From ToF-SIMS, we observe variations in the carbon, hydrogen, oxygen, chlorine, and fluorine content within the metal films and from nearby exposed substrate surfaces from qubit to qubit. We do not observe meaningful differences in the level of alloying at metal/metal or metal/substrate interfaces across qubits. Low temperature AFM reveals the presence of non-superconducting niobium hydrides at temperature $<$200K [24] and the density of these precipitates varies across qubits.
In terms of physical property measurements, DC transport measurements reveal variations in the normal-state resistance as well as switching current across the Josephson junction from qubit to qubit. From MO imaging, we observe variations in the total area of the region penetrated by magnetic flux as well as the non-uniformity of the flux front across qubits. And finally, from THz spectroscopy, we observe variations in the peak near-field THz signal measured at the edge of the superconducting pads from qubit-to-qubit. These observations were subsequently compared to the performance metrics of these qubits to identify properties and features in these devices that appear to trend with qubit quality factor.
Figure 2: A summary of the structural, chemical, and physical property differences from qubit-to-qubit within a single chip that were identified in this study. The technique through which each feature or property was measured is provided in the parenthetical.
II.2Relationship between Trench Depth and Performance
In agreement with previous reports by Gambetta et al. and Murray, we observe that there appears to be a relationship between increasing trench depth, which represents the depth of recesses etched into the underlying dielectric substrate immediately adjacent to superconducting metal edges, and increased qubit quality factor within qubits on a single chip in this study (Figure 3) [25, 18, 26]. An example of the relationship between trench depth and the quality factor of the qubit is provided in Figure S4 for superconducting qubits prepared with a niobium base metal capped with tantalum metal [1] on a silicon substrate.
@@
Based on electromagnetic (EM) simulations of the qubit geometry employed in this study, we observe that the energy participation ratio (EPR) of the lossy surface dielectric decreases logarithmically with increases in trench depth and saturates at trench depth values of the order of 70 nm, as seen in Figure 3b. We find that this reduction in the surface EPR primarily arises from a suppression in the field at the sidewalls (side EPR) of the superconducting metallization layer with increasing trench depth as opposed to suppression in the field from the top surface (top EPR). Qualitatively, these simulations agree with the experimental results regardless of the superconducting materials and substrates used, where, similarly, small variations in the trench depth trend with significant variations in the qubit quality factor for qubits with trench depth values less than 20 nm. With increasing trench depth, variations in trench depth appear to induce smaller variations in the qubit quality factor, and in the case where trench depths exceed 70 nm, trench depth appears to be uncorrelated with the qubit quality factor. However, these variations in trench depth do not entirely explain the qubit-to-qubit variations in performance observed in this study. Based on the simulated data, the shaded regions in 3a represent the level of qubit-to-qubit variation for a single chip expected in the case where trench depth represents the sole source of variation. This assumes a 10% error in the loss predicted by the simulations. Given that the performance of several qubits lies outside the shaded regions, it appears that other sources of variation contribute to the variations in performance observed. In the next sections, we will discuss two additional sources that correlate with performance. These sources are most apparent for qubits with substrate trench depths exceeding 70 nm (Chips 4 and 6) and qubits with a top surface consisting of niobium oxide (Chip 2).
Figure 5: EM simulations of the qubit sidewall. The sidewall angle ($\alpha$), oxide thickness (dh), radius of curvature of the top edge (R1), radius of curvature of the bottom edge (R2), distance of undercut ($x$), and undercut angle ($\beta$) are defined in (a). Surface EPR values are provided as a function of (b) $\alpha$ for dh = 5 nm and d = 100 nm as well as (c) dh for $\alpha$ = 0 degrees and d = 100 nm. Top represents the EPR associated with the top surface, side represents the EPR associated with the sidewall surface, and sum represents the total surface EPR, i.e. the sum of these two quantities.
II.3Sidewall Geometry
@@
As seen in Figure 5, we find a relationship between variations in $\alpha$ and the surface EPR of the sidewall. The simulations indicate that devices with $\alpha$ of approximately 30 degrees (as seen in Figure 4) exhibit surface EPR values that are 15-20% larger than a comparable device with $\alpha$ of approximately 10-15 degrees. On the other hand, we find that variations in the radius of curvature of the top edge have virtually no impact on the surface EPR due to competing effects between fields in the top surface and sidewall. Similarly, virtually no impact on surface EPR is observed when the radius of curvature of the base of the film varies, which also supports our experimental findings (Figure S6).
II.4Niobium oxide
@@
In particular, the data set associated with Chip 2 in Figure 3 represents a set of 4 qubits where the base metal consisted of bare niobium. As seen in Figure 6, although there is no correlation between trench depth and variations in qubit quality factor between devices, increases in niobium oxide thickness, specifically at the top surface, trend with decreases in quality factor. This agrees with the simulations in Figure 5. The bars in the scatter plot represent the range in thicknesses observed for each device and indicate that the surface oxide thickness of this polycrystalline film of niobium varies from 5 nm to 7 nm across a single 7.5 mm by 7.5 mm chip. The shaded regions in 6e represent the level of qubit-to-qubit variation expected in the case where the thickness of the niobium oxide represents the sole source of variation based on the simulated data.
@@
II.6Inconclusive Impact on Performance
@@
II.6Macroscopic Debris and Lithographic Defects
@@
II.6Josephson Junction: Structure and Electrical Properties
@@
As illustrated in Figures S9 and S10 and discussed previously, variations in grain structure and sidewall geometry of the Josephson Junction electrodes, as well as I-V response, are observed for the measured qubit devices. Concerning grain structure, we observe that the median grain size in the upper aluminum lead can vary from 30 nm to 50 nm, depending on the qubit. Similarly, in the lower aluminum lead, we observe that the median grain size varies from 20 nm to 35 nm. In both cases, there does not appear to be any relationship with the qubit quality factor. Similar variations to those observed in the sidewall angle of superconducting pads are observed for the aluminum leads, as captured in Figure S9. In this case, however, no clear trends between sidewall angle and qubit performance are observed. Finally, in terms of the I-V response, device-to-device variations are also observed in terms of normal state resistance as well as switching current (Figure S10). Once again, no trend is observed between these parameters and qubit performance, which thereby suggests that in the case of the high-coherence superconducting qubits investigated in this study, variations in the Josephson Junction morphology do not appreciably alter the overall performance.
II.6Surface Chemistry
@@
The resulting average impurity counts measured with ToF-SIMS are provided in Figure S11 Across the three analyzed chips, there do not appear to be any obvious trends between the surface chemistry of the superconducting pads and the qubit performance. Potentially, there may be a trend between residual fluorine content from plasma etching processes and performance, however, increased statistics are needed to isolate the impact of this impurity on performance. Moreover, although hydrogen impurities are known to bind with niobium at cryogenic temperatures and form non-superconducting precipitates, as reported by Lee et al. and seen in Figure S12, no consistent trend is observed between hydrogen impurity content and performance across all qubits[24]. This is in agreement with recent work suggesting that hydrogen impurities in niobium may not be linked to TLS loss [39]. Additionally, we also probed the chemistry of the substrate surface adjacent to the measured qubits. In this case, the spot-to-spot variation in the substrate surface chemistry for a single qubit was found to be similar to the variation in substrate surface chemistry observed across devices. Thus, dedicated studies where the surface chemistry is locally altered are needed to understand the impact of impurities in superconducting pads and substrate regions on device performance.
II.6Sidewall Oxide Thickness
@@
Similar to variations in the niobium top surface oxide thickness observed in Figure 6, variations in the niobium sidewall oxide thickness are also observed and reported. Unlike our observations for the niobium surface oxide, there does not exist an obvious correlation between variations in niobium sidewall oxide thickness and variations in performance. This may be due to two factors. First, unlike the 1-2 nm variations observed in the niobium surface oxide thickness, the variations in the niobium sidewall oxide thickness are of a smaller magnitude and typically range from 0.5 nm -1 nm. Once again, an additional dedicated study focused on varying the thickness of the niobium oxide at the sidewall is needed to meaningfully extract the relative impact variations in this parameter imposed on performance.
@@ III   Discussion
Concerning sidewall geometries, reproducible sidewalls can be achieved by developing more anisotropic dry etching recipes for the metal. To this end, sidewall imaging with electron microscopy can be a valuable tool for recipe optimization, while parameters such as etching chemistry (F-based vs Cl-based), level of physical bombardment (Ar content), and power are varied. Finally, with respect to surface encapsulation [1], our results indicate that incomplete coverage, especially close to the sidewall, can have a noticeable detrimental effect on device performance. This was the case for niobium films capped with AuPd as opposed to niobium films capped with Ta.